OpenStudy (anonymous):

From a spherical balloon of radius r, gas is escaping at the rate of 500cm^3/s. Find how fast the surface area A is shrinking at the instant r=250cm.

OpenStudy (anonymous):

let V be the volume then V=4/3pi r^3 dV/dt=4 pi r^2dr/dt=500cm^/s dr/dt=500/4 pi r^2 let A bethe area = 4pi r^2 dA/dt=8 pi r dr/dt = 8 pi r * 500 /4 pi r^2 = 1000/r when r = 250 dA/dt= 4cm^2/s

OpenStudy (amistre64):

From a spherical balloon of radius r, gas is escaping at the rate of 500cm^3/s. Find how fast the surface area A is shrinking at the instant r=250cm. \[\frac{dr}{dt}=\frac{dr}{dV}\frac{dV}{dt}\ ;\ \frac{dV}{dt}=500\] \[V = \frac{4}{3}pi\ r^3\ ;\ derive\ with\ repect\ to\ "V"\] \[\frac{dV}{dV} = \frac{4(3)}{3}pi\ r^2*\frac{dr}{dV}\] \[1 = 4pi\ r^2*\frac{dr}{dV}\] \[\frac{1}{4pi\ r^2}=\frac{dr}{dV}\] find the value of "r" and use it to determine dr/dV

OpenStudy (amistre64):

-500 might be more accurate, but just keep in mind that the radius is "shrinking" and apply the right sign afterwards :)

OpenStudy (amistre64):

..... and now that it reads surface area ... doh!!

OpenStudy (amistre64):

im going for a coffee :)